# Jack of All Trades, Master of Some, a Multi-Purpose Transformer Agent

### v2

@@ 2.2 Multi-Modal Transformer
While all the aforementioned approaches produce text as output and were solely trained to solve vision-and-language tasks such as Visual Q&A, our work spans both vision-and-language and embodied decision-making (e.g. control). Closer to our approach, a recent series of multi-modal Transformers trained for decision-making appeared. For instance, Jiang et al. (2022) trained a Transformer-based robot agent with Imitation Learning (IL) on tasks where a robot arm must place objects on a table to match an instruction. Their model uses multi-modal prompts as input (allowing them to define various objectives with a unified approach e.g. visual goals, demonstrations, text instructions, object naming) and directly produces motor actions for the robot arm. Similarly, RT-1 (Brohan et al., 2023b) leverages a multi-modal Transformer and IL to produce motor actions for robots but this time using expert demonstrations coming from real-world robots. Palm-E (Driess et al., 2023) and later RT-2 (Brohan et al., 2023a) proposed to study the impact of pretrained Transformers when learning decision-making robotics tasks with IL. Both works proposed to leverage a Visual Language Model (VLM) trained on vision-and-language tasks and finetune it for robotics tasks. While RT-2, being an extension of RT-1, directly outputs motor actions, Palm-E still produces, as its backbone VLM, sequences of text instructions (i.e. plans) that are then executed by low-level control policies.
@@ 3.2 Datasets
In this work, we have collected a wide range of datasets, classified into two main groups: sequential decision-making datasets and textual datasets. The former include a series of interaction sequences, each consisting of observations, actions and a subsequent rewards, generated by so-called expert agents, details of which are given in the following sections. The latter includes large corpora of textual data and image-text pairs. In order to promote the emerging field of general-purpose AI models, we have made these datasets, together with the expert agents and the full set of code required to generate them, available to the public as open resources in our Hugging Face repository, accessible at the following address https://huggingface.co/jat-project. To the best of our knowledge, this compilation is unprecedented in terms of the variety of tasks and the volume of data, representing a valuable new contribution to the field.
3.2.1 Sequential Decision-Making Datasets
Atari
BabyAI
Meta-World
MuJoCo
3.2.2 Text-Centric Datasets
Oscar
Conceptual-Captions
OK-VQA
Wikipedia
Refer to caption Refer to caption Refer to caption
@@ 3.3.2 Task-Specific Weight Adjustments
Sample Weight
Some tasks required more updates for effective convergence. To allow proportionate progress of all tasks during learning, these tasks are sampled more frequently. Specifically, Oscar, Conceptual-Captions and Wikipedia was assigned a sample weight of 10.0 while other have a sample weight of 1.0.
Loss Weight
Some control tasks require increased accuracy of actions. To allow for more strongly penalizing the error for these tasks, we assigned loss weights. In MuJoCo tasks, the loss weight is typically set at 10.0, except for the Pendulum task (20.0) and the Double Pendulum task (50.0). In Meta-World tasks, a uniform loss weight of 50.0 is used.
@@ Acknowledgements
CO2 Emission Related to Experiments
Experiments on the Jean Zay supercomputer (France), with a carbon efficiency of 57.3 gCO2eq/kWh (Statista Search Department, 2020), consumed 2000 GPU hours on Tesla V100-SXM2-32GB hardware (TDP [^5] Estimate for 11/2020: http://www.idris.fr/media/jean-zay/jean-zay-conso-heure-calcul.pdf: 482 W). Total emissions were estimated at 57.2 kgCO2eq, or 29.7 kgCO2eq when considering heat recovery. Estimations used the MachineLearning Impact calculator from (Lacoste et al., 2019).
@@ Appendix B JAT Dataset In Depth
This appendix completes the presentation of the JAT dataset begun in Section 3.2. This dataset, the scripts and expert agent used to generate it are available as open-source [^7] https://huggingface.co/jat-project/jat-dataset.
The JAT dataset contains sequential decision-making tasks from 4 domains: Atari, BabyAI, Meta-World and MuJoCo. Each sample in this dataset is an episode. This episode consists of a list of observations, actions and rewards, the nature and size of which depend on the task. In Figure 11, we represent for each Atari game the average return of the episodes of the dataset normalized by the human score from (Mnih et al., 2015). Notably, for 43 games the average score is higher than the human score, and for 31 games the average score is more than twice the human score. It should be noted, however, that for 7 games (Bowling, Montezuma’s Revenge, PitFall, Private Eye, Seaquest, Solaris and Venture) the average score is less than 10% of the human score.

### v3

@@ 2.2 Multi-Modal Transformer
These models, typically generating text outputs, are trained for vision-and-language tasks like Visual Q&A. However, recent multi-modal Transformers focus on decision-making. For example, Jiang et al. (2022) trained a robot with Imitation Learning (IL) using multi-modal prompts to produce motor actions. RT-1 (Brohan et al., 2023b) and RT-2 (Brohan et al., 2023a) use expert demonstrations for real-world robots, with RT-2 building on RT-1 by directly outputting motor actions. Palm-E (Driess et al., 2023) leverages a pretrained Visual Language Model (VLM) for robotics tasks, producing sequences of text instructions executed by control policies.
@@ 3.2 Datasets
In this work, we have collected a wide range of datasets, classified into two main groups: sequential decision-making datasets and textual datasets. The former include a series of interaction sequences, each consisting of observations, actions and a subsequent rewards, generated by so-called expert agents, details of which are given in the Appendix B. The latter includes large corpora of textual data and image-text pairs. In order to promote the emerging field of general-purpose AI models, we have made these datasets, together with the expert agents and the full set of code required to generate them, available to the public as open resources in our Hugging Face repository, accessible at the following URL https://huggingface.co/jat-project. To the best of our knowledge, this compilation is unprecedented in terms of the variety of tasks and the volume of data, representing a valuable new contribution to the field.
@@ 3.3.2 Task-Specific Weight Adjustments
Sample Weight Some tasks required more updates for effective convergence. To allow proportionate progress of all tasks during learning, these tasks are sampled more frequently. Specifically, Oscar, Conceptual-Captions and Wikipedia was assigned a sample weight of 10.0 while other have a sample weight of 1.0.
Loss Weight Some control tasks require increased accuracy of actions. To allow for more strongly penalizing the error for these tasks, we assigned loss weights. In MuJoCo tasks, the loss weight is typically set at 10.0, except for the Pendulum task (20.0) and the Double Pendulum task (50.0). In Meta-World tasks, a uniform loss weight of 50.0 is used.
@@ 4.1 Text-Centric Tasks
Refer to caption Refer to caption Refer to caption
@@ Appendix B JAT Dataset In Depth
B.1 Sequential Decision-Making Datasets
Atari
BabyAI
Meta-World
MuJoCo
Each sample in this dataset is an episode. This episode consists of a list of observations, actions and rewards, the nature and size of which depend on the task. In Figure 11, we represent for each Atari game the average return of the episodes of the dataset normalized by the human score from (Mnih et al., 2015). Notably, for 43 games the average score is higher than the human score, and for 31 games the average score is more than twice the human score. It should be noted, however, that for 7 games (Bowling, Montezuma’s Revenge, PitFall, Private Eye, Seaquest, Solaris and Venture) the average score is less than 10% of the human score.
@@ B.1 Sequential Decision-Making Datasets
B.1.1 Text-Centric Datasets
Oscar
Conceptual-Captions
OK-VQA
Wikipedia